# Quenched large deviations of Birkhoff sums along random quantum measurements

### v1

@@ 4.4 Regularity of the limiting objects
Fix $\alpha\in{\mathbb{R}}$ and let $\epsilon>0$. By Lemmas 4.3 and 4.8,
$\mathop{\mathrm{\mathstrut d}}\nolimits(Z_{0}^{(\alpha)},Z_{0}^{(\beta)})\leq\frac{1}{\delta}\left\|Z_{0}^{(\alpha)}-Z_{0}^{(\beta)}\right\|_{1}=\frac{1}{\delta}\left\|\Phi_{-N}^{(\alpha)}\cdot Z_{-N}^{(\alpha)}-\Phi_{-N}^{(\beta)}\cdot Z_{-N}^{(\beta)}\right\|_{1}.$
@@
$\lim_{\beta\to\alpha}\left\|\Phi_{-N}^{(\alpha)}\cdot Z_{-N}^{(\beta)}-\Phi_{-N}^{(\beta)}\cdot Z_{-N}^{(\beta)}\right\|_{1}=0.$ (4.21)
@@
$\mathrm{e}^{-C|\beta-\alpha|}\Phi_{-N}^{(\alpha)}\cdot Z\leq\Phi_{-N}^{(\beta)}\cdot Z\leq\mathrm{e}^{C|\beta-\alpha|}\Phi_{-N}^{(\alpha)}\cdot Z.$
@@
$\left\|\Phi_{-N}^{(\alpha)}\cdot Z_{-N}^{(\beta)}-\Phi_{-N}^{(\beta)}\cdot Z_{-N}^{(\beta)}\right\|_{1}\leq\sup_{Z\in\mathbb{S}}\left\|\Phi_{-N}^{(\alpha)}\cdot Z\right\|_{1}\left(\mathrm{e}^{C|\beta-\alpha|}-1\right).$
@@
$\lim_{\beta\to\alpha}\frac{1}{\beta-\alpha}\log\frac{\operatorname{tr}[\tilde{Z}_{1}^{(\alpha)}\varphi_{0}^{(\beta)}(Z_{-1}^{(\beta)})]}{\operatorname{tr}[\tilde{Z}_{1}^{(\alpha)}\varphi_{0}^{(\alpha)}(Z_{-1}^{(\beta)})]}=\frac{\operatorname{tr}\left[\tilde{Z}_{1}^{(\alpha)}\left.\frac{\mathop{\mathrm{\mathstrut d}}\nolimits\!missing}{\mathop{\mathrm{\mathstrut d}}\nolimits\!\gamma}\varphi_{0}^{(\gamma)}\right|_{\gamma=\alpha}(Z_{-1}^{(\alpha)})\right]}{\operatorname{tr}\left[\tilde{Z}_{1}^{(\alpha)}\varphi_{0}^{(\gamma)}(Z_{-1}^{(\alpha)})\right]}.\qed$
@@ 5 Proofs for Section 3
$=\operatorname{tr}(\tau Y\tau\varphi_{\omega}^{(1-\alpha)}(X)).$
@@ A.1 Almost sure convergence of convex functions
$\limsup_{N\to\infty}G_{N;\omega}(\alpha)\leq\limsup_{N\to\infty}\left(\frac{r-\alpha}{q-r}G_{N;\omega}(q)+\frac{\alpha-q}{q-r}G_{N;\omega}(r)\right)=\frac{r-\alpha}{q-r}\lambda(q)+\frac{\alpha-q}{q-r}\lambda(r).$
@@ A.2 A sub-additive ergodic upper bound for continuous families
$\limsup_{n\to\infty}\frac{1}{n}\log c_{n}^{(\alpha)}(\omega)\leq\inf_{n\to\infty}\frac{1}{n}\int\log c_{n}^{(\alpha)}\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}$ (A.7)
@@
$\begin{split}\limsup_{n\to\infty}\frac{1}{n}\log c_{n}^{(\alpha)}(\omega)&\leq\frac{1}{k}\int\log\left(\max\left\{t,c_{k}^{(\beta)}\right\}\right)\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}+\frac{1}{tk}\int\epsilon^{(k)}(\,\cdot\,,\delta)\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}\\ &\leq\frac{1}{k}\int\log\left(\max\left\{t,c_{k}^{(\alpha)}\right\}\right)\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}+\frac{2}{tk}\int\epsilon^{(k)}(\,\cdot\,,\delta)\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}.\end{split}$ (A.14)

### v2

@@ 4.4 Regularity of the limiting objects
Fix $\alpha\in{\mathbb{R}}$ and let $\epsilon>0$. By Lemmas 4.4 and 4.9,
$\mathop{\mathrm{\mathstrut d}}\nolimits(Z_{0}^{(\alpha)},Z_{0}^{(\beta)})\leq\frac{1}{\delta}\left\|Z_{0}^{(\alpha)}-Z_{0}^{(\beta)}\right\|_{1}=\frac{1}{\delta}\left\|\Phi_{-N}^{(\alpha)}\cdot Z_{-N}^{(\alpha)}-\Phi_{-N}^{(\beta)}\cdot Z_{-N}^{(\beta)}\right\|_{1},$
@@
$\lim_{\beta\to\alpha}\left\|\Phi_{-N}^{(\alpha)}\cdot Z_{-N}^{(\beta)}-\Phi_{-N}^{(\beta)}\cdot Z_{-N}^{(\beta)}\right\|_{1}=0,$ (4.25)
@@
$\mathrm{e}^{-C|\beta-\alpha|}\Phi_{-N}^{(\alpha)}\cdot Z\leq\Phi_{-N}^{(\beta)}\cdot Z\leq\mathrm{e}^{C|\beta-\alpha|}\Phi_{-N}^{(\alpha)}\cdot Z$
@@
$\left\|\Phi_{-N}^{(\alpha)}\cdot Z_{-N}^{(\beta)}-\Phi_{-N}^{(\beta)}\cdot Z_{-N}^{(\beta)}\right\|_{1}\leq\sup_{Z\in\mathbb{S}}\left\|\Phi_{-N}^{(\alpha)}\cdot Z\right\|_{1}\left(\mathrm{e}^{C|\beta-\alpha|}-1\right),$
@@
$\lim_{\beta\to\alpha}\frac{1}{\beta-\alpha}\log\frac{\operatorname{tr}[\tilde{Z}_{1}^{(\alpha)}\varphi_{0}^{(\beta)}(Z_{-1}^{(\beta)})]}{\operatorname{tr}[\tilde{Z}_{1}^{(\alpha)}\varphi_{0}^{(\alpha)}(Z_{-1}^{(\beta)})]}=\frac{\operatorname{tr}\left[\tilde{Z}_{1}^{(\alpha)}\left.\frac{\mathop{\mathrm{\mathstrut d}}\nolimits\!missing}{\mathop{\mathrm{\mathstrut d}}\nolimits\!\gamma}\varphi_{0}^{(\gamma)}\right|_{\gamma=\alpha}(Z_{-1}^{(\alpha)})\right]}{\operatorname{tr}\left[\tilde{Z}_{1}^{(\alpha)}\varphi_{0}^{(\alpha)}(Z_{-1}^{(\alpha)})\right]}$
by Proposition 4.10. ∎
@@ 5 Proofs for Section 3
$=\operatorname{tr}(\tau Y\tau\varphi_{\omega}^{(1-\alpha)}(X)),$
@@ A.1 Almost sure convergence of convex functions
$\limsup_{N\to\infty}G_{N;\omega}(\alpha)\leq\limsup_{N\to\infty}\left(\frac{r-\alpha}{r-q}G_{N;\omega}(q)+\frac{\alpha-q}{r-q}G_{N;\omega}(r)\right)=\frac{r-\alpha}{r-q}\lambda(q)+\frac{\alpha-q}{r-q}\lambda(r).$
@@ A.2 A sub-additive ergodic upper bound for continuous families
$\limsup_{n\to\infty}\frac{1}{n}\log c_{n}^{(\alpha)}(\omega)\leq\inf_{n\in{\mathbb{N}}}\frac{1}{n}\int\log c_{n}^{(\alpha)}\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}$ (A.7)
@@
$\begin{split}\limsup_{n\to\infty}\frac{1}{n}\log c_{n}^{(\alpha)}(\omega)&\leq\frac{1}{k}\int\log\left(\max\left\{t,c_{k}^{(\beta)}\right\}\right)\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}+\frac{1}{tk}\int\epsilon^{(k)}(\,\cdot\,,\delta)\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}\\ &\leq\frac{1}{k}\int\log\left(\max\left\{t,c_{k}^{(\alpha)}\right\}\right)\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}+\frac{2}{tk}\int\epsilon^{(k)}(\,\cdot\,,\delta)\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P},\end{split}$ (A.14)